OpenStudy (anonymous):

I will give you a medal Please name the 4 arithmetic operations that are performed on the functions. (please do NOT name the types of function but instead the operations of functions)

OpenStudy (zzr0ck3r):

Given the two functions \(f(x) \text{ and }g(x)\) we have \((f+g)(x)=f(x)+g(x)\\(f-g)(x)=f(x)-g(x)\\(fg)(x)=f(x)g(x)\\(\frac{f}{g})(x)=\frac{f(x)}{g(x)}\)

OpenStudy (anonymous):

okay thank you